# Beginner’s Charm: Beginner-Heavy Teams Are Associated With High Scientific Disruption

### v1

@@ Results
Importantly, teams with more beginners (Q4) maintain positive disruption–citation correlations for the broadest top-disruption range, whereas beginner-sparse teams (Q1) exhibit the fastest drop-off into negative territory (Figure 5). These quartile differences are robust across all cumulative top-$N$ thresholds: pairwise Fisher $r$-to-$z$ tests (Holm-adjusted) reveal significant divergence in $5$ of $6$ quartile pairs at every threshold ($P<10^{-300}$). The most pronounced gap appears between Q1 and Q4, maximizing at $N=100\%$ with a Fisher-transformed $z$-score difference of $\Delta z=-165.97$ ($P<10^{-300}$). Together, these findings suggest that while disruptive work is generally harder to recognize, highly disruptive papers from beginner-heavy teams are well-rewarded with high citation impact. This supports the notion that newcomers can succeed by taking bold intellectual leaps.
Discussion
@@
Teams are today the dominant engine of science, spanning nearly every field of inquiry and steadily replacing the lone researcher in driving knowledge forward [63, 55]. A central challenge in the science of science is to uncover what kinds of team dynamics foster originality and breakthroughs. Here we provide evidence for a universal and previously undocumented pattern: the presence of beginners—authors with no prior publication history—systematically enhances scientific disruption and innovation. This finding adds a critical new dimension to our understanding of how knowledge advances, challenging the prevailing assumption that beginners are merely blank slates or passive contributors.
@@
Our insights were made possible by a large-scale analysis of $28$ million papers across five decades. We find that beginner-heavy teams consistently produce more disruptive and innovative work across fields, time periods, and team sizes. This success is linked to distinctive knowledge-integration behaviors: beginners are less tethered to prevailing trends, more likely to pursue neglected or atypical references, and thus better positioned to recombine knowledge in novel ways. These mechanisms align with longstanding theories of creativity that emphasize originality through unusual combinations and broader search [64, 65, 14, 66, 67, 68, 69]. While philosophers of science describe knowledge growth as an endogenous process fueled by accumulated understanding, it requires engagement with a wide range of prior work—a condition beginners appear to fulfill despite having less time to absorb the field’s expanding burden of knowledge [14]. Unburdened by entrenched norms, beginners may be more open to exploring conceptual recombinations that others overlook. Their lower productivity relative to later-career researchers may also allow for more focused intellectual risk-taking, aligning with creativity theories that emphasize deep attention and cognitive agility as drivers of novel contributions [56, 46, 45].
@@
Collaboration dynamics also play a crucial role in shaping beginner success. We find that early-career collaborators are associated with greater disruption in beginner-heavy teams than senior ones, suggesting that younger collaborators may offer more compatible forms of intellectual flexibility and support. We additionally find that beginner-heavy teams thrive in disruptiveness when paired with collaborators who themselves have disruptive track records. These results contribute to a growing literature emphasizing how team composition, career age, diversity, and freshness [43, 34, 70, 33, 17, 71, 72, 73, 74, 75] mediate collaborative innovation. Encouragingly, we find that the highly disruptive works by beginner-heavy teams are rewarded with high citation impact in the academic literature, corroborating and further elucidating prior findings [61].
@@
Our work provides deeper nuances to contextualize previously reported insights on team disruption. For instance, prior work has shown that larger teams tend to be less disruptive than small ones [43], that team performance follows a U-shaped relationship with average career age [33], and that “fresh” teams (i.e., those without prior collaborations among members) are more likely to produce original or multidisciplinary work [34], among others. However, the role of absolute beginners with no prior publications has remained largely unexamined. To that end, we identify a pathway through which the well-known penalty of large team size on disruptiveness can be negated: beginner-heavy large teams can rival their smaller counterparts in disruptive capacity. Moreover, our results offer a finer-grained understanding of the left arm of the U-shaped curve relating average career age and disruption of teams [33], revealing the previously undocumented importance of the presence of absolute beginners in the low average career age territory. This is especially relevant given that over $98.2\%$ of scientific teams fall on the left side of that U-curve, underscoring the widespread applicability of our findings. We also extend earlier findings about team freshness by showing that not only do fresh collaborations enhance disruption by bringing in diverse expertise to the team, but so too does individual-level freshness, who bring in no track record. This complements and deepens prior understandings of team novelty [34], offering a more granular account of how new entrants inject creative potential into scientific teams.
@@
Our study has limitations. We define beginners solely by career age, without distinguishing differences in physical age or the amount of prior training. We also adopt a team-level lens, without tracking individual beginners longitudinally to evaluate whether their disruptive success persists or peaks later in their careers [76, 59, 77, 78, 79]. Future work should explore these trajectories, as well as the micro-dynamics of collaboration: how beginner ideas are introduced, evaluated, and either nurtured or sidelined. Moreover, examining whether similar patterns hold in domains such as startups, patents, or the arts could reveal whether the “beginner’s charm” generalizes beyond academia. While the disruption and atypical combination scores we used capture fundamentally different dynamics—paradigm-shifting effects in downstream literature versus novelty in recombining prior literature, respectively [80, 44, 81, 42]—future work can apply additional indicators of innovation, including natural language processing-based measures to analyze abstract-level novelty, conceptual distance, or idea emergence over time. Importantly, our findings are limited to correlational observations rather than making causal claims.
@@
The implications of these findings are wide-ranging and challenge the prevailing emphasis on experience and track record in science hiring, funding, and team assembly. First, they offer guidance for team science: teams designed for disruptive outcomes may benefit from intentionally integrating beginners alongside collaborators with disruptive track records. This contrasts sharply with current industry trends that prioritize track record over untapped potential—exemplified by Meta’s aggressive push to “build the most elite and talent-dense team in the industry” [82]. Second, at the policy level, funding agencies could create programs that explicitly support beginner-led projects or incentivize collaborations that amplify beginner contributions. Such efforts would counteract systemic biases that currently privilege safe, incremental research over risky exploration [83, 20]. Finally, at the institutional level, universities and research organizations might reconsider evaluation systems overly tied to citation counts and productivity (i.e., the so-called ‘publish or perish’ culture [84]), shifting toward measures that reward originality, risk-taking, and long-term vision [85].
@@
Overall, our results highlight a form of “beginner’s charm” in science: newcomer-heavy teams occupy a unique position to disrupt, driven by their fresh perspectives and willingness to explore less-traveled intellectual paths. As science becomes more collaborative, and as the academic workforce ages [86, 14, 87, 88, 89], the under-tapped value of newcomers becomes increasingly vital to shape the future trajectory of discovery.
Materials and Methods
Dataset
@@
In this paper, we analyze SciSciNet V2, a refreshed update to SciSciNet. It is a large-scale, integrated data lake derived from the Microsoft Academic Graph (MAG) [90] and OpenAlex [91] to support science of science research. The dataset contains publication information on 100,418,971 authors and 249,803,279 publication records. The author names are disambiguated in the database [37]. We select all articles published between 1971 and 2020 across 172 fields, leading to a total of 35,751,547 articles. Furthermore, we remove extreme entries (authors with $1000$ or more overall publications, authors whose career span is $80$ or more, and beginner authors with $10$ or more papers in their debut year) to curate a final dataset of 28,042,889 papers for this analysis. We explain the dataset preprocessing steps in the Supplementary Text.
Disruption Score
@@
Citation networks are a standard tool for measuring scientific disruption—identifying works that introduce new directions rather than build incrementally on prior research. The Disruption Score $D$ quantifies this by comparing how future papers cite the focal paper and its references [42, 43]:
@@
Here, $N_{i}$ is the number of papers that cite the focal paper but not its references, $N_{j}$ cite both, and $N_{k}$ cite only the references. The score ranges from 1-1 (developmental) to +1+1 (disruptive). Positive values suggest the focal paper has shifted the research trajectory, while negative values indicate work that deepens existing lines. We use the Disruption Score provided in the SciSciNet dataset [35].
Atypical Combination Score
Building on the view that scientific advances arise from recombining existing knowledge [14, 53, 54, 55, 92], we quantify atypicality following Uzzi et al. [53]. For each focal paper, all unordered pairs of reference venues in its reference list are formed, and the typicality of each pair is assessed using the literature preceding the paper’s publication year. Let $O_{ij}$ denote the observed prior co‐occurrence count of venues $i$ and $j$. $O_{ij}$ is compared to an empirical null generated by Monte Carlo edge‐switching on the global citation network that preserves (i) each paper’s in‐ and out‐degree and (ii) the forward/backward temporal distribution of citations, thereby holding constant the propensity of papers to cite and be cited over time. For each pair $(i,j)$, a null mean $\mu_{ij}$ and standard deviation $\sigma_{ij}$ are obtained across randomized networks, and a normalized score is computed:
@@
Here, positive $z_{ij}$ values indicate conventional venue pairings (more frequent than expected by chance), whereas negative $z_{ij}$ values indicate atypical venue pairings (less frequent than expected). A paper’s overall atypicality is summarized by the median of its pairwise zz–scores (median zz[53]. Lower median $z$ values correspond to more atypical (novel) combinations; higher values indicate more conventional combinations. In this paper, we use the median $z$ values for papers provided by the SciSciNet data lake [35], which computes zz–scores using the procedure above with the prior‐year literature as the reference set. We restrict analyses to papers with at least two cited venues (so that at least one pair exists). We report the median $z$ as the “Atypical Combination Score,” with lower scores indicating greater novelty.
@@
Ethics Approval This research utilizes public data and therefore does not require IRB approval.
@@
Data, Materials, and Software Availability The data used in this paper are publicly available at https://northwestern-cssi.github.io/sciscinet/. Our end-to-end data preprocessing and statistical analysis code can be found at https://github.com/cssai-research/beginners_charm.
@@
Supplementary Materials for Beginner’s Charm: Beginner-Heavy Teams Are Associated With High Scientific Disruption
Dataset
In this paper, we analyze SciSciNet V2, a refreshed update to SciSciNet. It is a large-scale, integrated data lake derived from the Microsoft Academic Graph (MAG) [90] and OpenAlex [91] to support science of science research. The dataset contains publication information on 100,418,971 authors and 249,803,279 publication records. The author names are disambiguated in the database [37]. We select all articles published between 1971 and 2020 across multiple fields, leading to a total of 35,751,547 articles. Furthermore, we remove extreme entries (authors with $1000$ or more overall publications, authors whose career span is $80$ or more, and beginner authors with $10$ or more papers in their debut year) to curate a final dataset of 28,042,889 papers for this analysis.
@@
We utilized BigQuery [93], a Google Cloud Platform enterprise-level data warehouse tool, to preprocess SciSciNet data for our research. We first added a debut_year column in the Authors table based on each author’s earliest publication (any doctype—that is, preprint, dissertation, article, review, book chapter, book, dataset, or retraction). After that, we constructed the All_Yearly_Author_Profiles table, where (authorid, year) is a composite key. For a given year $y$, this table contains the career_age of authorid in year $y$, the number_of_papers by authorid in year $y$, and aggregated metrics (productivity, average citation, average disruption, etc.) regarding their previous year’s publications. Then, we enhanced the Paper table by adding pre-publication factors needed for our research. For each year $y$ ranging from 1971197120202020; and for each paperid published in year $y$, we looked up the rows for (authorid, $y$) to construct team-level pre-publication factors of a paper such as beginner_author_count, beginner_author_ratio, average_disruption_of_team, and average_citation_count_of_team, by aggregating the retrieved rows for the (authorid, $y$) key for each author of the paperid. We discarded records where doi or disruption was missing. Furthermore, we removed some unrealistic entries (authors with $1000$ or more overall publications, authors whose career age is $80$ or more, and beginner authors with $10$ or more papers in their debut year), which constituted 19.16519.165% of the data. However, all of our findings remain true even if we include these unrealistic entries. Finally, we used the rank function of Pandas [94] to calculate the disruption and citation percentiles of each paper.
@@
Figure S2: Beginner author ratio positively correlates with higher disruption percentiles across all decades. This figure shows the relationship between the beginner author ratio and disruption percentile, broken down by team size and publication decade (1971–2020). Across all decades and team sizes, a higher share of beginner authors is associated with more disruptive work.
Figure S3: Beginner author ratio positively correlates with higher disruption percentiles across different disciplines. This figure shows the relationship between the beginner author ratio and disruption percentile, broken down by discipline and team size. In nearly all fields, increasing the share of beginner authors correlates with higher disruption.
Figure S4: Beginner author ratio negatively correlates with atypical combination score percentiles (i.e., positively correlates with innovation) across all team configurations. This figure demonstrates the negative correlation between the proportion of beginner authors and atypical combination score percentile across varying team sizes.
Figure S5: Beginner authors tend to cite less popular papers. This figure displays the ratio of beginner authors and their average reference popularity percentiles. Beginner authors tend to cite less popular papers.
@@
8+ 0.991 [0.940, 1.000]*** 0.855 [0.489, 1.000]*** -1.000 [-1.000, -1.000]***
Table S2: Correlation between beginner author ratio and disruption percentile across decades in the overall dataset. All correlations are statistically significant, demonstrating the robustness of this association over time.
Group NN Pearson $r$ $P$ Spearman $\rho$ $P$ Kendall $\tau$ $P$
All Data 28,042,889 0.105901 0.0 0.081011 0.0 0.062118 0.0
1971-1980 1,477,772 0.068475 0.0 0.053241 0.0 0.040983 0.0
1981-1990 2,418,668 0.073201 0.0 0.052333 0.0 0.040204 0.0
1991-2000 3,757,769 0.075007 0.0 0.050995 0.0 0.038976 0.0
2001-2010 7,290,555 0.095750 0.0 0.067300 0.0 0.051500 0.0
2011-2020 13,098,125 0.136875 0.0 0.109910 0.0 0.084223 0.0
Table S3: Correlations between author beginner ratio and disruption percentile by field in the overall dataset. Pearson’s $r$, Spearman’s $\rho$, and Kendall’s $\tau$ with significance stars: $^{***}P<0.001$.

### v2

@@ 2.3 Highly disruptive papers by beginner-heavy teams are highly cited
Importantly, teams with more beginners (Q4) maintain positive disruption–citation correlations for the broadest top-disruption range, whereas beginner-sparse teams (Q1) exhibit the fastest drop-off into negative territory (Figure 5). These quartile differences are robust across all cumulative top-$N$ thresholds: pairwise Fisher $r$-to-$z$ tests (Holm-adjusted) reveal significant divergence between every quartile pair at every threshold ($P<10^{-300}$). The most pronounced gap appears between Q1 and Q4, maximizing at $N=100\%$ with a Fisher-transformed $z$-score difference of $\Delta z=-185.91$ ($P<10^{-300}$). Together, these findings suggest that while disruptive work is generally harder to recognize, highly disruptive papers from beginner-heavy teams are well-rewarded with high citation impact. This supports the notion that newcomers can succeed by taking bold intellectual leaps.
2.4 Robustness to alternative career-stage definitions
As a robustness check, we refine our career-stage definitions by splitting early-career authors into early (1–5 years) and mid-career (6–10 years) groups, corresponding roughly to graduate training versus early faculty stages, while keeping the beginner and senior definitions unchanged. The beginner category (0 years) is fixed by construction, as it captures authors with no prior publications. The senior threshold (11+ years) reflects a decade or more of publishing, a period reasonably sufficient for disciplinary enculturation and the accumulation of domain-specific expertise. We therefore focus the refinement on the intermediate career range, where heterogeneity in training stage and professional autonomy is most pronounced.
This alternative specification yields the same qualitative patterns across the main results. In the overall dataset, the beginner ratio remains positively associated with disruption (Pearson’s $r=0.11$), the early-career ratio is much weaker ($r=0.005$), the mid-career ratio is slightly negative (r=0.002r=-0.002), and the senior ratio remains most negative (r=0.08r=-0.08), preserving the same graded decline across career stages from beginners to seniors. The team-size-stratified analyses likewise preserve the ordinal pattern: beginner ratios remain most positively associated with disruption across team sizes, early- and mid-career ratios show weaker intermediate associations, and senior ratios remain negative (Supplementary Figure S11 and Supplementary Table S8).
The collaborator-composition analyses are similarly stable under the four-way partition: the disruption heatmaps continue to show that teams become more disruptive as the beginner ratio rises, while the corresponding atypical-combination heatmaps preserve the same directional ordering of innovation-related patterns across collaborator mixes (Supplementary Figures S12S13). Taken together, these checks show that our conclusions do not depend on treating years 1–10 as a single class, but instead reflect a stable career-stage gradient in which beginner-heavy teams remain the most disruptive and innovative.
3 Discussion
@@
Teams are today the dominant engine of science, spanning nearly every field of inquiry and steadily replacing the lone researcher in driving knowledge forward [wuchty2007increasing, jones2008multi]. A central challenge in the science of science is to uncover what kinds of team dynamics foster originality and breakthroughs. Here we provide evidence for a near-universal and previously undocumented pattern: the presence of beginners—authors with no prior publication history—systematically enhances scientific disruption and innovation. This finding adds a critical new dimension to our understanding of how knowledge advances, challenging the prevailing assumption that beginners are merely blank slates or passive contributors.
@@
Our insights were made possible by a large-scale analysis of $29$ million papers across eight decades. We find that beginner-heavy teams consistently produce more disruptive and innovative work across disciplines, time periods, team sizes, and alternative definitions of career stages. This success is linked to distinctive knowledge-integration behaviors: beginners are less tethered to prevailing trends, more likely to pursue neglected or atypical references, and thus better positioned to recombine knowledge in novel ways. These mechanisms align with longstanding theories of creativity that emphasize originality through unusual combinations and broader search [uzzi2005collaboration, schilling2005small, jones2009burden, schoenmakers2010technological, kelley2013breakthroughs, didegah2013factors, mammola2021impact]. While philosophers of science describe knowledge growth as an endogenous process fueled by accumulated understanding, it requires engagement with a wide range of prior work—a condition beginners appear to fulfill despite having less time to absorb the field’s expanding burden of knowledge [jones2009burden]. Unburdened by entrenched norms, beginners may be more open to exploring conceptual recombinations that others overlook. Their lower productivity relative to later-career researchers may also allow for more focused intellectual risk-taking, aligning with creativity theories that emphasize deep attention and cognitive agility as drivers of novel contributions [li2024productive, park2023papers, chu2021slowed].
@@
Collaboration dynamics also play a crucial role in shaping beginner success. We find that early-career collaborators are associated with greater disruption in beginner-heavy teams than senior ones, suggesting that younger collaborators may offer more compatible forms of intellectual flexibility and support. We additionally find that beginner-heavy teams thrive in disruptiveness when paired with collaborators who themselves have disruptive track records. These results contribute to a growing literature emphasizing how team composition, career age, diversity, and freshness [wu2019large, zeng2021fresh, hofstra2020diversity, yang2024unveiling, packalen2019age, kwiek2024young, grant1996toward, kunze2011age, schneid2016age, li2021leveraging] mediate collaborative innovation. Encouragingly, we find that the highly disruptive works by beginner-heavy teams are rewarded with high citation impact in the academic literature, corroborating and further elucidating prior findings [li2024breaking].
@@
Our work provides deeper nuances to contextualize previously reported insights on team disruption. For instance, prior work has shown that larger teams tend to be less disruptive than small ones [wu2019large], that team performance follows a U-shaped relationship with average career age [yang2024unveiling], and that “fresh” teams (i.e., those without prior collaborations among members) are more likely to produce original or multidisciplinary work [zeng2021fresh], among others. However, the role of absolute beginners with no prior publications has remained largely unexamined. To that end, we identify a pathway through which the well-known penalty of large team size on disruptiveness can be negated: beginner-heavy large teams can rival their smaller counterparts in disruptive capacity. Moreover, our results offer a finer-grained understanding of the left arm of the U-shaped curve relating average career age and disruption of teams [yang2024unveiling], revealing the previously undocumented importance of the presence of absolute beginners in the low average career age territory. This is especially relevant given that over $98.2\%$ of scientific teams fall on the left side of that U-curve, underscoring the widespread applicability of our findings. We also extend earlier findings about team freshness by showing that not only do fresh collaborations enhance disruption by bringing in diverse expertise to the team, but so too does individual-level freshness, who bring in no track record. This complements and deepens prior understandings of team novelty [zeng2021fresh], offering a more granular account of how new entrants inject creative potential into scientific teams.
@@
Our study has limitations. We define beginners solely by career age, without distinguishing differences in physical age or the amount of prior training. We also adopt a team-level lens, without tracking individual beginners longitudinally to evaluate whether their disruptive success persists or peaks later in their careers [liu2018hot, li2025quantifying, wang2019early, yin2019quantifying, liu2021understanding]. Future work should explore these trajectories, as well as the micro-dynamics of collaboration: how beginner ideas are introduced, evaluated, and either nurtured or sidelined. Moreover, examining whether similar patterns hold in domains such as startups, patents, or the arts could reveal whether the “beginner’s charm” generalizes beyond academia. While the disruption and atypical combination scores we used capture fundamentally different dynamics—paradigm-shifting effects in downstream literature versus novelty in recombining prior literature, respectively [wang2023effect, leibel2024we, leahey2023types, funk2017dynamic]—future work can apply additional indicators of innovation, including natural language processing-based measures to analyze abstract-level novelty, conceptual distance, or idea emergence over time. Importantly, our findings are limited to correlational observations rather than making causal claims.
@@
The implications of these findings are wide-ranging and challenge the prevailing emphasis on experience and track record in science hiring, funding, and team assembly. First, they offer guidance for team science: teams designed for disruptive outcomes may benefit from intentionally integrating beginners alongside collaborators with disruptive track records. This stands in tension with two ongoing trends. On one hand, industry hiring is increasingly privileging elite track records—exemplified by Meta’s recent push to “build the most elite and talent-dense team in the industry” [duffy2025meta]. On the other, AI is disproportionately displacing entry-level roles, reducing traditional pathways through which newcomers gain experience and contribute intellectually [wef2026_entry_level_ai]. Together, these forces risk systematically excluding precisely the beginner contributions that our results identify as critical for innovation. Second, at the institutional level, universities and research organizations might reconsider evaluation systems overly tied to citation counts and productivity (i.e., the so-called ‘publish or perish’ culture [plume2014publish]), shifting toward measures that reward originality, risk-taking, and long-term vision [reisz2022loss]. Finally, at the policy level, funding agencies could create programs that explicitly support beginner-led projects or incentivize collaborations that amplify beginner contributions. Such efforts would counteract systemic biases that currently privilege safe, incremental research over risky exploration [rzhetsky2015choosing, livan2019don].
@@
Overall, our results highlight a form of “beginner’s charm” in science: newcomer-heavy teams occupy a unique position to disrupt, driven by their fresh perspectives and willingness to explore less-traveled intellectual paths. As science becomes more collaborative, and as the academic workforce ages [milojevic2018changing, jones2009burden, matthews2011aging, sugimoto2016age, blau2017us], the under-tapped value of newcomers becomes increasingly vital to shape the future trajectory of discovery.
4 Materials and Methods
4.1 Dataset
@@
In this paper, we analyze SciSciNet V2, a refreshed update to SciSciNet. It is a large-scale, integrated data lake derived from the Microsoft Academic Graph (MAG) [wang2020microsoft] and OpenAlex [openalex] to support science of science research. The dataset contains information on 249,803,279 publication records. The author names are disambiguated in the database [torvik2009author]. We select all articles published between 1941 and 2020 across all 19 Level-0 (top-level) disciplines. Furthermore, we remove extreme entries (authors with $1000$ or more overall publications, authors whose career span is $80$ or more, authors who publish more than $50$ papers in a year, and beginner authors with $10$ or more papers in their debut year) to curate a final dataset of 29,054,261 papers for this analysis. However, all of our findings remain true even if we include these extreme entries. Please note that we do not include earlier decades due to a lack of samples available (e.g., after filtration, only 2037 papers remain between 1931-1940, and even fewer for earlier decades). We explain the dataset preprocessing steps in the Supplementary Text.
4.2 Disruption Score
@@
Citation networks are a standard tool for measuring scientific disruption—identifying works that introduce new directions rather than build incrementally on prior research. The Disruption Score $D$ quantifies this by comparing how future papers cite the focal paper and its references [funk2017dynamic, wu2019large]:
@@
Here, $N_{i}$ is the number of papers that cite the focal paper but not its references, $N_{j}$ cite both, and $N_{k}$ cite only the references. The score ranges from 1-1 (developmental) to +1+1 (disruptive). Positive values suggest the focal paper has shifted the research trajectory, while negative values indicate work that deepens existing lines. We use the Disruption Score provided in the SciSciNet dataset [lin2023sciscinet].
4.3 Atypical Combination Score
Building on the view that scientific advances arise from recombining existing knowledge [jones2009burden, uzzi2013atypical, guimera2005team, jones2008multi, yang2022gender], we quantify atypicality following Uzzi et al. [uzzi2013atypical]. For each focal paper, all unordered pairs of reference venues in its reference list are formed, and the typicality of each pair is assessed using the literature preceding the paper’s publication year. Let $O_{ij}$ denote the observed prior co‐occurrence count of venues $i$ and $j$. $O_{ij}$ is compared to an empirical null generated by Monte Carlo edge‐switching on the global citation network that preserves (i) each paper’s in‐ and out‐degree and (ii) the forward/backward temporal distribution of citations, thereby holding constant the propensity of papers to cite and be cited over time. For each pair $(i,j)$, a null mean $\mu_{ij}$ and standard deviation $\sigma_{ij}$ are obtained across randomized networks, and a normalized score is computed:
@@ 4.3 Atypical Combination Score
Here, positive $z_{ij}$ values indicate conventional venue pairings (more frequent than expected by chance), whereas negative $z_{ij}$ values indicate atypical venue pairings (less frequent than expected). A paper’s overall atypicality is summarized by the median of its pairwise zz–scores (median zz[uzzi2013atypical]. Lower median $z$ values correspond to more atypical (novel) combinations; higher values indicate more conventional combinations. In this paper, we use the median $z$ values for papers provided by the SciSciNet data lake [lin2023sciscinet], which computes zz–scores using the procedure above with the prior‐year literature as the reference set. We restrict analyses to papers with at least two cited venues (so that at least one pair exists). We report the median $z$ as the “Atypical Combination Score,” with lower scores indicating greater novelty.
@@
Ethics Approval This work uses public data and does not require IRB approval.
@@
Data and Code Availability The data is publicly available at https://northwestern-cssi.github.io/sciscinet/. Our end-to-end preprocessing and statistical analysis code is available at https://github.com/cssai-research/beginners_charm.
@@
Supplementary Information for Beginner’s Charm: Beginner-Heavy Teams Are Associated With High Scientific Disruption
@@
We utilized BigQuery [google2025bigquery], a Google Cloud Platform enterprise-level data warehouse tool, to preprocess SciSciNet data for our research. We first added a debut_year column in the Authors table based on each author’s earliest publication. After that, we constructed the All_Yearly_Author_Profiles table, where (authorid, year) is a composite key. For a given year $y$, this table contains the career_age of authorid in year $y$, the number_of_papers by authorid in year $y$, and aggregated metrics (productivity, average citation, average disruption, etc.) regarding their previous year’s publications. Then, we enhanced the Paper table by adding pre-publication factors needed for our research. For each year $y$ ranging from 1941194120202020; and for each paperid published in year $y$, we looked up the rows for (authorid, $y$) to construct team-level pre-publication factors of a paper such as beginner_author_count, beginner_author_ratio, average_disruption_of_team, and average_citation_count_of_team, by aggregating the retrieved rows for the (authorid, $y$) key for each author of the paperid. We discarded records where doi or disruption was missing. Furthermore, we removed some unrealistic entries (described in the main manuscript). Finally, we used the rank function of Pandas [mckinney2010pandas] to calculate the disruption and citation percentiles of each paper.
@@
Figure S2: Beginner author ratio positively correlates with higher disruption percentiles across all decades. This figure shows the relationship between the beginner author ratio and disruption percentile, broken down by team size and publication decade (1941–2020). We plot scatter points with at least $1000$ papers. Across all decades and team sizes, a higher share of beginner authors is associated with more disruptive work.
Figure S3: Beginner author ratio positively correlates with higher disruption percentiles across all 19 broad disciplines. This figure shows the relationship between the beginner author ratio and disruption percentile, broken down by discipline and team size. Across disciplines, increasing the share of beginner authors correlates with higher disruption.
Figure S4: Beginner author ratio negatively correlates with atypical combination score percentiles (i.e., positively correlates with innovation) across team sizes. This figure demonstrates the negative correlation between the proportion of beginner authors and atypical combination score percentile across varying team sizes.
Figure S5: Beginner author ratio negatively correlates with average reference popularity across team sizes. This figure displays the ratio of beginner authors and their average reference popularity percentiles. Teams with a higher beginner ratio tend to cite less popular papers.
@@
Figure S11: Beginner-heavy teams are disruptive across team sizes, robustly for a 4-way career stage split. We show the relationship between disruption percentiles and author ratios at different career stages, split by team sizes. Beginner author ratios are positively associated with higher levels of disruption, while teams with higher senior author ratios tend to produce less disruptive work. The results are robust to team size. Shaded regions denote 9595% C.I.
Figure S12: Early-career collaborators are associated with higher disruption in beginner-heavy teams under a four-way career-age split. We illustrate disruptions of pairwise combinations of beginner, early-career, mid-career, and senior authors across different author ratios. The heatmap cells visualize median disruption percentiles, with darker shades indicating higher disruption. While teams heavier on beginners are progressively more disruptive, early-career collaborators are associated with better beginner disruption than mid-career or senior collaborators.
Figure S13: Early-career collaborators are associated with higher innovation in beginner-heavy teams under a four-way career-age split. We illustrate atypical combination scores of pairwise combinations of beginner, early-career, mid-career, and senior authors across different author ratios. The heatmap cells visualize median atypical combination score percentiles, with lighter shades indicating higher innovation. While teams heavier on beginners are progressively more innovative, early-career collaborators are associated with better beginner innovation than mid-career or senior collaborators.
@@
8+ 0.931 [0.738, 1.000]*** 0.848 [0.517, 1.000]*** -0.939 [-1.000, -0.714]***
Table S2: Correlations between beginner author ratio and disruption percentile across decades in the overall dataset. All reported correlations are statistically significant at $P<0.001$ (denoted by ***), demonstrating the robustness of this association over time.